OpenStudy (anonymous):

Solve the equation p - 4 = 10 for p. Please show your work and solution in the space provided. How the heck....

OpenStudy (lilg132):

p -4 = 10 10 + 4 = p 14 = p

OpenStudy (anonymous):

7/8*p-4=10 sorry didnt write it out properly.

OpenStudy (lilg132):

7/8p - 4 = 10 7/8p = 10 + 4 7/8p = 14 7/8p / (7/8) = 14 / (7/8) p = 16

OpenStudy (lilg132):

i find it easier turning the fractions into numbers and then working it out so 7/8 = 0.875 so the equation would be 0.875p - 4 = 10 and then you rearrange it bringing 4 to the other side (remember it wil change its value e.g. -4 bringint it to the other side will give you + 4) 0.875p = 14 and now to get p on its own you need to divide by the number next to it and you also divide the other side by the same value so you would get 0.875p / 0.875 = 14 / 0.875 which gives you p = 16

OpenStudy (anonymous):

thank you

OpenStudy (lilg132):

your welcome